### v1

@@ Abstract
Vertical Federated Learning (VFL) is an emergent distributed machine learning paradigm wherein owners of disjoint features of a common set of entities collaborate to learn a global model without sharing data. In VFL, a host client owns data labels for each entity and learns a final representation based on intermediate local representations from all guest clients. Therefore, the host is a single point of failure and label feedback can be used by malicious guest clients to infer private features. Requiring all participants to remain active and trustworthy throughout the entire training process is generally impractical and altogether infeasible outside of controlled environments. We propose Decoupled VFL (DVFL), a blockwise learning approach to VFL. By training each model on its own objective, DVFL allows for decentralized aggregation and isolation between feature learning and label supervision. With these properties, DVFL is fault tolerant and secure. We implement DVFL to train split neural networks and show that model performance is comparable to VFL on a variety of classification datasets.
@@ 1 Introduction
Federated Learning McMahan et al. (2017), or FL, was introduced by Google researchers as a strategy for distributed learning, addressing communication efficiency and data privacy. Distributed participants in FL training do not expose their data to any other party. Instead, they train local models on their data, guided by a coordinator that has no knowledge of the agents’ data. Thus, FL extends collaborative machine learning to contexts where data communication is undesirable or forbidden. Recent advances in communication systems and cloud computing infrastructure combined with an increased public awareness and legislation regarding data privacy (such as the EU’s GDPR) have contributed to FL’s advancement as an emergent machine learning paradigm.
@@
Among FL’s developments is Vertical Federated Learning (VFL). Standard FL can be said to have a “horizontal” or “sample-parallel” division of data, i.e. each participant holds unique samples within a shared feature space. On the other hand, VFL participants hold unique features of a common sample space, which is a “vertical” division, and is “feature-parallel”. This allows guest agents who individually have incomplete information about their target to learn meaningful joint representations without sharing data. A typical example: an imaging center, pathology lab and general hospital may come together to predict rare types of cancer, even though health regulations such as HIPAA may disallow them from collecting each others’ data.
VFL guest agents undertake an “entity alignment” phase before training, where they ensure their datasets’ horizontal indexing is identical, i.e., all participants process features belonging to the same entity in any round of training. At each round, each guest passes its features to its local “guest model.” Guests share their local representations with a host. For example, a guest training a vertically federated ANN may send its activations. A boosted decision tree guest may send a weak learner. The host owns data labels and is aware of the identity of the sample or batch that the guests are processing at any time. The host also possesses a “host model” capable of aggregating these representations and making predictions against the labels. The host calculates a loss based on its aggregation and provides feedback (e.g. gradients) to the guests, which then begin the next round of training. We refer to the combined system of guest and host models as VFL models.
VFL has found increasing success in recent years, with various real-world implementations and proposals Liu et al. (2023). However, VFL suffers from some flaws which make its implementation impractical where participants are few in number and each computationally powerful (as in cross-silo scenarios), and entirely infeasible otherwise. One such flaw is that the link between the guests and the host is a single point of failure. Training a batch of data is contingent on the exchange of intermediate results between the guests and the host. If the host waits indefinitely for all guests to communicate, a guest crashing or losing connection would completely halt training. If the host sets a deadline, some inputs to its model will remain unfulfilled. (In the latter case, a late message will also cause a catastrophic fault). If gradients from the host to the guests are lost or arrive in the incorrect order, guest model convergence is hampered. One might dismiss such events as infrequent or at least controllable in cross-silo settings. However, that is naive. When some of the guests (and even the host) could be geographically sparse and/or low-power devices, it is catastrophic.
A fundamental limitation of VFL is imposed by the size of the joint sample space. For instance, databases in groups of large nationwide financial institutions may share many but not all customers. Guests must discard data outside the intersection, which could have been used to learn better guest models. In systems where the intersection is small, there may not be enough samples to train a VFL model altogether. In general, this also hinders VFL’s applicability to systems with large-scale participation. Consider again the hospital example. Suppose the task is now to predict bone marrow cancer. Along with the pathology lab, the imaging center and the general hospital, we also include data collected by an orthopaedic hospital. Even if the overlap of patients at the other facilities is significant, their overlap with patients of the orthopaedic hospital may be too small.
Another drawback of VFL (and FL in general) is the risk of inference attacks from curious guests. Intuitively, a host makes predictions based on information it receives from all guests. The updates or feedback it returns is derived from this prediction and the ground truth. Therefore, feedback sent to one guest contains some mutual information with respect to the others and to the label. Luo et al. (2021); Fu et al. (2022) show that a malicious agent can engineer its messages to the host such that the label and the other guests’ features can be conveniently extracted from the gradients it receives in return. Attempts to safeguard VFL against inference attacks are generally based on differential privacy, homomorphic encryption, and multi-party techniques such as audits and shared secrets Yang et al. (2023); Khan et al. (2023). These operations incur a high computational cost, which might not be feasible outside some cross-silo scenarios. Further, these methods hamper model convergence.
Building intricate mechanisms to mitigate these issues inflates the complexity of the problem. In this work, we attack the source directly. Backpropagating gradients to guests allows inference attacks. The wait for a response or feedback causes synchronization issues. What if we eschewed end-to-end backpropagation (BP) altogether? We present Decoupled Vertical Federated Learning (DVFL), a novel strategy for ANNs to address these shortcomings. The underlying principle is blockwise learning, where gradient-isolated groups of layers of an ANN are trained on local objectives. Such greedily trained ANNs have been shown to perform comparably to those trained by BP Belilovsky et al. (2019); Journé et al. (2023). We decouple guest models from host models by training them on their own local objectives. Similarly, we decouple supervision from aggregation by training host models on local objectives. We use a final transfer learning model to make predictions against labels. This yields two useful system properties. One, this allows redundant hosts. Standard VFL would require hosts to share labels with each other, which is forbidden. Two, we may use separate datasets for feature extraction and label inference.
With redundancy, even if a guest’s connection to one host fails, training may continue as normal on others. If hosts may fail, the redundancy also affords some degree of tolerance. By using separate datasets for inference and representation learning, we may train guest and host models on data which is not necessarily labeled and not necessarily a part of a joint sample space. This makes DVFL applicable to systems with small intersections of sample spaces. Further, the label owner’s supervised model can be trained completely offline, without the risk of faults.
In Section 2, we provide a formal description of VFL and some algorithms which implement it for various machine learning models. We also elaborate on the shortcomings we hope to address with DVFL. In Section 3, we review some closely-related lines of research and use them to motivate DVFL. In Section 4, we provide a formal description of DVFL with an implementation for ANNs. We show some baseline results on benchmark machine learning tasks in Section 5. In the Appendix, we present more results and discuss our design choices. The main contributions of this work are:
@@
    We present DVFL, a strategy which eschews host-to-guest feedback as well as label-to-host feedback and therefore has inbuilt redundancy, security and fault handling (including faults during the backward pass) while performing comparably to VFL.
@@ 2 Background: Standard VFL
Entities $\mathbf{x}$ are realizations of a fixed distribution $X$ (say). The rows of a dataset are features of a unique entity $\mathbf{x}$. We may serialize these entities using a sample index. Let us say the set of sample indices corresponding to entities the host has labels for is $\mathcal{S}_{\mathrm{labels}}$. Let us say the entity with index $j$ is $\mathbf{x}_{j}$.
@@ 2.1 The VFL Hierarchy
Guests: $\mathcal{G}$ is the set of feature-owning collaborators: the guests. ||$\cdot$|| ⋅ | denotes cardinality. Each $g_{i}\in\mathcal{G}$ has its own
  1. 1.
    Dataset $\mathcal{D}_{i}=\left\{\mathbf{x}_{j,\mathcal{F}_{i}}\in\mathbb{R}^{|\mathcal{F}_{i}|}:j\in\mathcal{S}_{i}\right\}$, where $\mathcal{F}_{i}\subseteq\{1,2,\cdots,D\}$ is the subset of $D$ features that the guest owns, and $\mathcal{S}_{i}$ is the set of entities $g_{i}$’s dataset has records for. $\mathbf{x}_{j,\mathcal{F}_{i}}$ are the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$.
  2. 2.
    Guest model $m_{i}(\text{ }\cdot\text{ };\theta_{i}):\mathbb{R}^{|\mathcal{F}_{i}|}\to\mathbb{R}^{k_{i}}$ with parameters $\theta_{i}$.
The guest model $m_{i}$’s task is to concisely represent the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$, for all sample indices $j$ that are in the intersection of all guests, i.e. $\forall j\in\mathcal{S}_{\mathrm{guests}}$, where $\mathcal{S}_{\mathrm{guests}}:=\bigcap_{i=1}^{|\mathcal{G}|}\mathcal{S}_{i}$ and pass them to the host $h$. Guests do not have access to other guests’ model parameters or data features. In general, they do not have access to labels, either.
Host: The host agent $h$ executes every round of training. The host has access to target labels. The host, given representations for all available features of $\mathbf{x}_{j}$, may use its own model to make useful predictions for $\mathbf{x}_{j}$. The host has
  1. 1.
    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
  2. 2.
    Host model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{h}$.
It is possible that $h\in\mathcal{G}$ i.e. the host may also be one of the guests. In such situations, the host is often known as the “active party” and guests as “passive parties.”
If the host has access to the full computational graph, it may adjust all parameters to better predict label $\mathbf{y}$, usually by minimizing the expected value of some loss function. Otherwise, it may update only its parameters $\theta_{h}$ and send intermediate results (e.g. gradients) via which the guests may update their own models. Together, the guests and host solve the following joint optimization problem:
@@
2.2 Private Set Intersection & Entity Alignment
All guests process entities in the same order, each passing the next entity to its local model. For this to happen, all guests must know the set of common entities $\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$. This is usually achieved using entity resolution (if required) and some form of Private Set Intersection (PSI).
PSI is a multiparty computation of the intersection of sets belonging to each party without revealing members outside the intersection, typically by exchanging encrypted versions of the set Freedman et al. (2004). Entity resolution is a multiparty computation to generate a set of unique identifiers from features which are not strictly unique or precise. Entity resolution may be performed by comparing cryptographic long keys Hardy et al. (2017); Nock et al. (2018). After the host obtains the results of PSI, there is a phase of entity alignment. In this phase, the host decides the order of entities which all guests must follow and informs them.
@@ 2.3 Split Training for Neural Networks
A faithful implementation of VFL for ANNs is SplitNN Gupta & Raskar (2018). SplitNN was originally proposed as an algorithm to distribute the training of a global neural network across two agents by splitting the network depth-wise into two. At the “split layer”, forward and backward signals are communicated between the two devices. [^1] One might say that the standard SplitNN addresses a special case of VFL where the host has access only to labels and the (single) guest has access only to features. The original paper Gupta & Raskar (2018) also presents a protocol for training when the features are horizontally distributed across $N$ guests. Ceballos et al. (2020); Romanini et al. (2021) extend SplitNN for vertically partitioned data by vertically splitting the shallower layers (see Figure 1).
@@
Effectively, the guest models $m_{i}$ are ANNs and host model $m_{h}$ consists of a concatenation layer $m_{h}^{0}$ followed by an ANN $m_{H}^{1}$.
@@ 2.4 Impracticalities in VFL Training
Vertical Federated Learning is contingent on implicit assumptions regarding the reliability, trustworthiness and scale of participation.
Subversion of these assumptions can be catastrophic. A lost guest output or gradient is a single point of failure. The intersection $\mathcal{S}_{\mathrm{labels}}\cap\mathcal{S}_{\mathrm{guests}}$ may not be meaningfully large, especially when there are many guests. And, it is possible for a malicious guest to infer private data from feedback.
Thus far, the application of VFL has been limited to a small number of cross-silo settings where the silo owners are powerful enough to assume reliability and to afford the overhead of privacy guarantees. Further, cross-silo participation is small in number so scalability is not a concern. VFL is altogether infeasible for a large class of systems where devices are many or have low compute and communication bandwidths.
@@ 2.4.1 Single Points of Failure
  • Guest faults: A guest device may be unable to compute its output, update its model, or communicate.
@@
From a VFL host’s perspective, a guest fault and communication fault manifest as a missing input to its model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ during training. The host may employ one of two strategies:
@@
    • Under a guest fault or communication fault, the host will remain waiting indefinitely, and training will not continue.
@@
    • Under a guest fault or communication fault, training will fail since the host model is missing input.
@@
Therefore, a communication or guest fault leads to a catastrophic failure for training. When the host fails, the guest models are unable to update their parameters. Since the cardinality of the joint training dataset is that of the intersection of participants’ sample spaces $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$ (which may be small), losing updates from a single batch of data may be detrimental to model performance.
@@ 2.4.3 Inference Attacks
Fu et al. (2022) demonstrate that it is possible for a malicious agent to directly infer the true class from gradients (e.g. by analyzing their sign) for many common loss functions during training. The authors also show that after VFL training is complete, it is possible for a curious guest to train its own prediction model if it can gain access to a very small number of labels. If a malicious guest inflates its learning rate during training, the aggregator gives more importance to results from its model. This way, a prediction model trained only on the features of the malicious guest is almost identical to the aggregator model itself. A malicious guest equipped with such a prediction head can make accurate guesses for labels without accessing other guests’ features.
@@
Typical defenses include differential privacy, gradient compression, homomorphic encryption and secret shares for gradients and model features. However, these methods usually have a negative effect on model convergence rate and tightness Khan et al. (2023). Moreover, implementing such cryptographic algorithms is computationally expensive, and the cost increases with the number of participants.
@@ 3 Related work
3.1 Fault-Tolerant Distributed Optimization
The impact of redundancy for fault tolerance in distributed systems is well studied. Liu et al. (2021) provide a rigorous analysis of redundancy in the case of distributed optimization, showing that systems which are $2f$-redundant are approximately resilient to $f$ Byzantine agents. The authors also present two fault-tolerant algorithms for horizontally distributed convex optimization. Liu (2021); Bouhata et al. (2022) extensively review approaches to fault tolerance in distributed learning. Most methods require synchronization by an aggregator and incur computational overhead. Further, most methods involve manipulating gradients, which forces a choice between convergence rate and tightness.
For VFL: Few methods exist for to address fault tolerance in VFL. Li et al. (2023) introduce FedVS, an algorithm for systems where guest models are polynomial networks Livni et al. (2014a; b). FedVS uses a form of secret sharing to introduce redundancy in the input data, and can tolerate a fixed number of late or failed guests. However, FedVS and many of its predecessors Shi et al. (2022); Chen et al. (2020a); Hu et al. (2019) only address faults in the forward pass.
3.2 Federated Transfer Learning
Federated Transfer Learning (FTL) is a strategy for systems where there is some overlap in both sample space and feature space. The host owns labels and some features (i.e. it is an active participant) for entities in its sample space. Guests have limited labels for their task, but have some small overlap in feature space and sample space with the host (“aligned samples”).
Liu et al. (2020) present Secure Transfer Federated Learning (STFL), a framework for a single host and a single guest. Both parties learn a common latent space using local feature extractors. Encodings of the passive participant’s unaligned samples in this feature space can be used by the passive participant to learn its task model. Feng & Yu (2020); Feng et al. (2022) extend this idea to multiple parties.
For the related case of VFL with small $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$, Sun et al. (2023) introduce few-shot Vertical Federated Learning, where a few epochs of training on the aligned dataset is followed by semi-supervised learning on the non-intersecting entities of each client’s dataset. To generate labels, guests cluster the gradients they receive from the aligned samples. While this enables remarkable model performance, there is unwanted computational overhead required to generate the labels. Moreover, this is a form of label inference attack! Not only is this a breach of privacy to some degree, this clustering could be used downstream by a malicious agent to perform feature inference attacks (see Section 2.4.3).
3.3 Greedy and Localized Learning
Algorithms which update parameters in hidden layers of ANNs based on local or auxiliary objectives (as opposed to the global label loss) have recently seen renewed interest. In the scope of VFL training, they enable some useful properties for ANN training.
By eschewing the wait to evaluate the global objective, most greedy algorithms are backward unlocked if not update unlocked. In order to partially address forward locking for model parallel training Belilovsky et al. (2020) propose a replay buffer mechanism. Each agent responsible for training a module uses a replay buffer to store activations from the previous module’s agent. During a forward pass, the agent reads an activation from the buffer. If the forward signal has propagated through the previous module, it would have written the latest activation to the buffer. If not, the previous activation is reused.

### v2

@@ Abstract
Vertical Federated Learning (VFL) is an emergent distributed machine learning paradigm for collaborative learning between clients who have disjoint features of common entities. However, standard VFL lacks fault tolerance, with each participant and connection being a single point of failure. Prior attempts to induce fault tolerance in VFL focus on the scenario of “straggling clients”, usually entailing that all messages eventually arrive or that there is an upper bound on the number of late messages. To handle the more general problem of arbitrary crashes, we propose Decoupled VFL (DVFL). To handle training with faults, DVFL decouples training between communication rounds using local unsupervised objectives. By further decoupling label supervision from aggregation, DVFL also enables redundant aggregators. As secondary benefits, DVFL can enhance data efficiency and provides immunity against gradient-based attacks. In this work, we implement DVFL for split neural networks with a self-supervised autoencoder loss. When there are faults, DVFL outperforms the best VFL-based alternative (97.58% vs 96.95% on an MNIST task). Even under perfect conditions, performance is comparable.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Federated Learning (McMahan et al., 2017), or FL, was introduced by Google researchers as a strategy for distributed learning, addressing communication efficiency and data privacy. Distributed participants in FL training do not expose their data to any other party. Instead, they train local models on their data, guided by a coordinator that has no knowledge of the agents’ data. Thus, FL extends collaborative machine learning to contexts where data communication is undesirable or forbidden. Recent advances in communication systems and cloud computing infrastructure combined with an increased public awareness and legislation regarding data privacy (such as the EU’s GDPR) have contributed to FL’s advancement as an emergent machine learning paradigm.
@@
Among FL’s developments is Vertical Federated Learning (VFL). Standard FL can be said to have a “horizontal” or “sample-parallel” division of data: each participant holds unique samples within a shared feature space. Conversely, VFL participants hold unique features of a common sample space, which is a “vertical”, “feature-parallel” division. This allows guest agents who individually have incomplete and information about their target to learn meaningful joint representations without sharing data.
VFL has received nascent interest, with some real-world proposals and applications (Liu et al., 2023). However, VFL requires careful synchronization which makes it difficult to engineer and scale. Critically, training is contingent on the exchange of intermediate results between the guests and the host. A crashed connection or participant on the backward pass means a missed model update. On the forward pass, it is catastrophic: an unfulfilled model input.
We propose Decoupled Vertical Federated Learning (DVFL), where guests and hosts each train asynchronously on their own unsupervised objective before the label owner learns a transfer learning model from its labels. At an algorithmic level, DVFL avoids the forward and backward locking between hosts and guests. Hence, it is free of a single point of failure. Such “localized” training means that other participants can continue training even when one fails. Further, separating label inference from aggregation allows for redundant aggregators, further localizing faults.
The exchange of intermediate results is also the root of many other limitations of VFL including inference attacks. Feedback received from the host has mutual information with labels and other guests’ features, which can be exploited by curious guests (mitigating this is typically costly in computation and performance: see Yang et al. (2023); Khan et al. (2023)). Another is the inability to learn from features pertaining to samples outside the intersection of all participants’ sample spaces. DVFL does not suffer from these problems.
Altogether, we expect DVFL to enable learning from vertically partitioned data in scenarios where they were so far infeasible, and help scale implementations where they weren’t. The main contributions of this work are:
@@
    We present DVFL, a strategy that eschews both host-to-guest and label-to-host feedbacks, and therefore enables redundancy, has inbuilt fault handling (including those during the backward pass) and inherent gradient privacy, and performs comparably to VFL.
@@ 2 Background: Standard VFL
2.1 Notation
The rows of a dataset are features of a unique entity $\mathbf{x}$. We may serialize these entities using a sample index e.g. as $\mathbf{x}_{j}$. $g_{i}$ is the $i^{\text{th}}$ data-owning agent, member of set of guests $\mathcal{G}$. $\mathcal{F}_{i}$ is the set of features that guest $g_{i}$ records. $\mathbf{x}_{j,\mathcal{F}_{i}}$ are the features $\mathcal{F}_{i}$ of the $j^{\text{th}}$ entity. $m_{i}(\text{ }\cdot\text{ };\theta_{i})$ is $g_{i}$’s model, $\mathbb{R}^{|\mathcal{F}_{i}|}\to\mathbb{R}^{k_{i}}$ function parameterized by $\theta_{i}$, which is learnable. $\mathcal{S}_{i}$ is the set of sample indices for which $g_{i}$ has records. $\mathcal{S}_{\text{guests}}$ = $\bigcap_{i=1}^{|\mathcal{G}|}$. $h$ is the host agent, which owns labels. $\mathcal{S}_{\text{labels}}$ is the set of sample indices corresponding to labeled entities. $m_{h}(\text{ }\cdot\text{ };\theta_{h})$ is the host’s $\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\dots\times\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\mathrm{out}}$ learnable model, which accepts the outputs of $m_{i}$ and makes a task-relevant prediction.
@@ 2.2 The VFL Hierarchy
Guest: The guest role does not entail access to labels or other guests’ data. The guest’s private model $m_{i}$’s task is to concisely represent the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$ $\forall$ sample indices $j\in\mathcal{S}_{\text{guests}}\cap\mathcal{S}_{\text{labels}}$. Guests obtain these encodings and pass them to the host $h$.
Host: The host agent $h$ executes every round of training. The host has access to target labels. The host, given representations for all available features of $\mathbf{x}_{j}$, may use its own model to make useful predictions for $\mathbf{x}_{j}$.
If the host has access to the full computational graph, it may adjust all parameters to better predict label $\mathbf{y}$, usually by minimizing the expected value of some loss function. Otherwise, it may update only its parameters $\theta_{h}$ and send intermediate results (e.g. gradients) via which the guests may update their own models.
Together, the guests and host solve the following joint optimization problem:
@@
It is possible that the host may also be one of the guests. In such situations, the host is often known as the “active party” and guests as “passive parties.”
@@ 2.3 Split Training for Neural Networks
SplitNN (Gupta & Raskar, 2018) is a faithful implementation of VFL for NNs. It was originally proposed as an algorithm to distribute the training of a global neural network across two agents by splitting the network depth-wise into two. At the “split layer”, forward and backward signals are communicated between the two devices. [^1] One might say that the standard SplitNN addresses a special case of VFL where the host has access only to labels and the (single) guest has access only to features. The original paper (Gupta & Raskar, 2018) also presents a protocol for training when the features are horizontally distributed across $N$ guests. Ceballos et al. (2020); Romanini et al. (2021) extend SplitNN for vertically partitioned data by vertically splitting the shallower layers (see Figure 1).
@@
Effectively, the guest models $m_{i}$ are NNs and host model $m_{h}$ consists of a concatenation layer $m_{h}^{0}$ followed by an NN $m_{h}^{1}$.
@@ 2.4 Impracticalities in VFL Training
VFL is contingent on implicit assumptions on the reliability, trustworthiness and scale of participation. As discussed in the introduction, subversion of these assumptions can be catastrophic. A lost guest output or gradient is a single point of failure. The intersection $\mathcal{S}_{\mathrm{labels}}\cap\mathcal{S}_{\mathrm{guests}}$ may not be meaningfully large, especially when there are many guests. And, it is possible for a malicious guest to infer private data from feedback. We further analyze these shortcomings here.
@@ 2.4.1 Single Points of Failure
  • Guest faults: A guest may be unable to compute its output, update its model, or communicate.
@@
From a VFL host’s perspective, a guest fault and connection fault manifest as a missing input to its model $m_{h}(\text{ }\cdot\text{ };\theta_{h})$ during training. The host may employ one of two strategies:
@@
    • Under a guest fault or connection fault, the host will remain waiting indefinitely, and training will not continue.
@@
    • Under a guest fault or connection fault, training will fail since the host model is missing input.
@@
Therefore, a connection or guest fault leads to a catastrophic failure for training. When the host fails, the guest models are unable to update their parameters. Since the cardinality of the joint training dataset is that of the intersection of participants’ sample spaces $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$ (which may be small), losing updates from a single batch of data may be detrimental to model performance.
@@ 2.4.3 Inference Attacks
Fu et al. (2022a) demonstrate that it is possible for a malicious agent to directly infer the true class from gradients (e.g. by analyzing their sign) for many common loss functions during training. If a curious guest inflates its learning rate, the joint model gives preference to its outputs. If it can obtain even a small number of labels, it may train its own prediction head. The authors show that such a malicious model can make meaningful guesses for labels.
@@
Typical defenses include differential privacy, gradient compression, homomorphic encryption and secret shares for gradients and model features. However, these methods usually have a negative effect on model convergence rate and tightness (Khan et al., 2023). Moreover, implementing such cryptographic algorithms is computationally expensive, and scales poorly.
@@ 3 Related Work
In this section, we review some closely-related lines of research which have motivated DVFL. A more verbose discussion is available in the Appendix E.
Fault Tolerant Distributed Optimization
Like any other distributed system, redundancy enables fault resilience and tolerance in distributed optimization (Liu et al., 2021). Liu (2021); Bouhata et al. (2022) provide extensive reviews of Byzantine fault tolerant horizontally distributed learning algorithms. Most methods hurt model performance and carry significant computational overhead.
For VFL: For VFL, there is limited work on crash faults. However, there is a line of work on the related case of straggling guests. That problem definition only extends to the count of failures being bounded as in FedVS (Li et al., 2023), and/or late messages eventually arriving (Xia et al., 2023; Shi et al., 2022; Zhang et al., 2021; Hu et al., 2019). A real system could have an arbitrary number and duration of crashes, so these methods do not generalize to crash faults. Typically, they do not address faults in the backward pass. Paradigms that require periodic synchronization e.g. Gradient Assisted Learning (Diao et al., 2022) are similarly hamstrung.
Asynchronous VFL
There are some efforts in speeding up training by relaxing synchronization in and splitting effort of gradient calculation. But many of these methods expect all inputs during the forward pass (Wang et al., 2024; Shi et al., 2024; Zhang et al., 2021; Gu et al., 2021). Chen et al. (2020a) present a VFL algorithm which has an asynchronous forward pass, which is a viable candidate for crash tolerance. Another line of work (Zhang et al., 2021; Gu et al., 2021) has an implicitly asynchronous forward pass– its aggregation strategy is summation rather than concatenation.
Federated Transfer Learning
Federated Transfer Learning (FTL) is a strategy for systems with overlap in both sample space and feature space. In Liu et al. (2020); Feng & Yu (2020); Feng et al. (2022), parties learn a common latent space using local feature extractors and transfer to their tasks. For vertically partitioned datasets with small $|\mathcal{S}_{\mathrm{labels}}|$ but large $|\mathcal{S}_{\mathrm{guests}}|$, Sun et al. (2023) introduce Few-Shot VFL. A few epochs of training on the aligned dataset are followed by clustering-based semi-supervised learning on the unlabelled entities of each client’s dataset. Besides the unwanted computational overhead required to generate the labels, this is effectively a label inference attack.
Greedy and Localized Learning
Recently, there has been a resurgence of interest in algorithms which train modules of NN layers on local objectives as opposed to BP. One reason is inherent locking in BP, described in Jaderberg et al. (2017). With local supervision, a layer need not wait for the global backward signal to update its model (“backward unlocking”). Belilovsky et al. (2020) propose a replay buffer: Each module of layers stores its inputs in a replay buffer and forward passes its front. It need not wait for a fresh forward signal to compute an output (“forward unlocking”). Note that forward and backward locking at the split layer are exactly the failure points for VFL. Local updates have proved useful for practical HFL (e.g. Collins et al. (2022); Jeon & Kim (2022); Briggs et al. (2020)), where labels are available to guests. For VFL, Fu et al. (2022b) use cached statistics to construct local gradients. As far as we know, DVFL is the first with fully localized model training.
4 Why Not an Engineering-Level Solution?
Keen observers will note that the critical failures of the “wait” and “timeout” strategies discussed in §2.4.1 can be partially mitigated with some extra bookkeeping. If the guest sends back the sample index (or indices of samples in its batch) and epoch number with each message, the host can uniquely identify which request it belongs to. Now timeout may be employed with a strategy to fill missing inputs; late messages may be identified and discarded. Wait may be employed by periodically repeating unfulfilled requests; messages from other iterations may be queued. So why DVFL?
Notably, neither strategy handles faults on on backward passes. Moreover, wait is a busy-wait. Besides being incompatible with deadline-based scheduling and wasting computational and communication bandwidth with repeated unserviced requests, there is no guarantee that training ever continues. In fact, this would also be the case for event-driven computation. But it is not so for DVFL. As for timeout, it is not trivial to choose a good replacement for missing inputs. We investigate some obvious options and show that DVFL outperforms them.